Average cost of 5 apples and 4 mangoes is Rs. 36. The average cost of 7 apples and 8 mangoes is Rs. 48. Find the total cost of 24 apples and 24 mangoes.

Grade:

1 Answers

avinash sharma
14 years ago

hi ..the answer of this question is 2088

Explanation:

Average cost of 5 apples and 4 mangoes = Rs. 36
Total cost = 36 * 9 = 324

Average cost of 7 apples and 8 mangoes = 48
Total cost = 48 * 15 = 720

Total cost of 12 apples and 12 mangoes = 324 + 720 = 1044
Therefore, cost of 24 apples and 24 mangoes = 1044 * 2 = 2088